### v1

# Ticket-based multi-strand method for increased efficiency in proof-of-work based blockchains

Abstract
This paper outlines a method aiming to increase the efficiency of proof-of-work based blockchains using a ticket-based approach. To avoid the limitation of serially adding one block at a time to a blockchain, multiple semi-independent chains are used such that several valid blocks can be added in parallel, when they are added to separate chains. Blocks are added to different chains, the chain index being determined by a “ticket” that the miner must produce before mining a new block. This allows increasing the transaction rate by several orders of magnitude while the system is still fully decentralized and permissionless, and maintaining security in the sense that a successful attack would require the attacker to control a significant portion of the whole network.
1 Introduction
A proof-of-work blockchain system as described in [1] provides several interesting features. In particular, such systems can be:
  • Decentralized: no central authority
  • Permissionless: anyone can decide to participate
Information is stored in a distributed ledger, copies of which can exist in many places but no instance has authority over the others. The proof-of-work mechanism provides a form of collective decision making regarding how information can be added to the distributed ledger.
One drawback of such blockchain systems is that the rate at which information can be added to the ledger is rather limited. This is because information is always added one block at a time, serially, at an even pace. For example, for the Bitcoin blockchain a new block is added about every 10 minutes, while for Monero the corresponding time is 2 minutes [2].
For cryptocurrency applications, each block contains one or more transactions, and one way to increase the rate of transactions is to increase the maximum allowed block size. However, the block size cannot be increased infinitely; the work involved in added a block needs to be feasible. In practice, even if the possibilities of increasing block size and reduced block time are taken into account, it appears unlikely that blockchains such as Bitcoin and Monero as currently designed could handle the transaction rates that would be necessary if such cryptocurrencies were to be used by a significant portion of the world’s population. That would require billions of transactions per day, several orders of magnitude more than what seems possible due to the current limitation of only adding one block at a time, serially.
In this paper we discuss a way to significantly increase the efficiency of proof-of-work blockchains by exploiting parallelism, while keeping the system decentralized and permissionless. The described method builds on similar ideas as discussed in [3] but realized in a different way. The improvement in throughput compared to a single-strand blockchain depends on the chosen amount of parallelism. As an example, a multi-stranded chain of 1024 strands would allow 1024 times as many transactions per day.
2 Method
In this section we first describe the standard (single-strand) proof-of-work blockchain as proposed in [1] for comparison, and then we describe our proposed multi-strand method.
2.1 Standard (single-strand) blockchain
In a proof-of-work blockchain as described in [1] each block contains the hash of the previous block, a set of transactions, and a nonce:
The work of producing a new valid block involves finding a nonce such that the hash of the new block begins with a certain number of zero bits.
When a new valid block has been found, that new block is sent to other nodes, and the other nodes verify that it is indeed a valid block by checking the transactions and the previous hash as well as the hash of the new block itself. When other nodes have verified that the new block is valid, each of them add it to their local chain. They have incentive to do so because the longest chain is considered correct, so everyone wants their new blocks to be added to a chain that is as long as possible.
2.2 Ticket-based multi-strand blockchain
When there is only one chain where a new block can be added, only one of the nodes (miners) that are trying to produce a new block can succeed, while all the other attempts to create new blocks will fail. One obvious way to allow more than one to succeed is to have several independent chains; if each chain can grow independently of the others, then several valid blocks, one per chain, can be produced in parallel. However, having completely independent chains would be bad for security since an attacker could pick one chain to attack and would only need to redo the proof-of-work for the selected chain in order to rewrite history for that chain.
We want to keep the property that an attacker needs to control a large part of the CPU power of the entire network, while we at the same allow parallel growth as if we had several independent chains. To achieve this, we devise a “ticket-based” approach as follows.
Instead of a single chain, we consider a different form of distributed ledger that consists of $n=2^{p}$ chains which we label “Chain 0” to “Chain $n-1$”. Each chain works similarly to the single-strand case, with the difference that each block is required to include a valid ticket:
The ticket is a data structure that contains the hash of the last known block from each of the $n$ chains, together with a nonce chosen such that the hash of the ticket satisfies the following: (1) the hash starts with a certain number of zero bits and (2) the last $p$ bits of the hash correspond to the chain index of the chain for which the ticket is valid.
The procedure to follow for a miner who wants to add a new block now consists of two proof-of-work phases, scanning for first a ticket nonce and then a block nonce, as follows:
  • First create a ticket by scanning for a ticket nonce that makes the ticket satisfy the criteria (1) and (2) above. Now the miner knows which chain it is possible to add a block to.
  • Then include that ticket in a new candidate block for the chain that the ticket is valid for, and scan for a block nonce that makes the block valid.
Note that the procedure above involves two different nonce values, a ticket nonce and a block nonce. In both cases the proof-of-work difficulty is determined by the number of zero bits required. The difficulties should be selected such that both parts require significant work.
Note also that the miner cannot choose which chain to add a block to, since the chain index is determined by the ticket. The miner tries to find a nonce leading to a hash starting with enough zero bits, while some other bits of the same hash determine the chain index. The chain index becomes effectively random and the ticket thus functions as a lottery ticket giving the miner access to mine a block for one specific chain. If a miner were to insist on only wanting to add blocks to one specific chain, then that miner would need to discard all tickets produced for other chains and the probability of getting a ticket with the desired chain index is only $1/n$ so the cost in wasted CPU power would be very high. Thus, miners have a strong incentive to let the bits of the ticket determine the chain to work on, leading to effectively random distribution among the $n$ chains. An attacker who wants to target a specific chain would still need to control a large part of the CPU power of the whole network.
The procedure to check if a new block is valid is the same as for the single-strand case except that the ticket is also verified, as follows:
  • Calculate the hash of the ticket and check the last $p$ bits to find the chain index.
  • Look at the information stored in the ticket about the hash of the last block for that chain index, checking that it in fact matches the hash of the last block for that chain.
  • Check that the required number of zero bits are found in the beginning of the hash of the ticket.
The purpose of requiring the ticket to contain the hash of the last block of each chain is to ensure that the ticket is fresh; without that requirement it would be possible for an attacker to spend a long time hoarding tickets to carry out an attack against a specific chain later. The requirement to include the hash of the last block of each chain in the ticket makes such hoarding impossible since old tickets will no longer be valid.
To illustrate how the multi-strand method gives increased efficiency, consider the following example: we take the simplest case $n=2$ so there are two chains, Chain 0 and Chain 1. Imagine there are 100 different miners. Each miner starts by creating a ticket, and they look at the last bit of the ticket hash to determine which chain they can mine on. About 50 miners will get Chain 0 and the other 50 will get Chain 1. One miner succeeds in finding a new block for Chain 0, and at about the same time another miner succeeds in finding a new block for Chain 1. Both of them distribute their new blocks to other miners, and both blocks are accepted because they are not in conflict, the two blocks are added to the two separate chains, in parallel. If a single-strand chain had been used then only one of the blocks could have been successfully added. The increased efficiency lies in the fact that two blocks are added to the distributed ledger instead of just one.
3 Discussion
The multi-strand approach discussed in this paper allows increased efficiency because new valid blocks can be added in parallel, which becomes possible when using a set of semi-independent chains instead of a single chain.
The idea of using several chains to increase efficiency is not new; one example of previous work employing this idea is the “Parallel Chains” work of Fitzi et al [3]. However, although a similar idea is involved, their method differs significantly from the ticket-based approach outlined in the present paper. In [3] the miner needs to prepare a metablock effectively containing candidate blocks for each of the $n$ chains, something that is not needed in our ticket-based approach which should therefore have better scalability for large values of $n$. The ticket-based method is also simpler in the sense that no special “synchronization chain” is involved, all $n$ chains are treated in the same way, none of them is special and no synchronization between chains is needed.
Regarding security, the worst case would be an attacker who wants to target one specific chain and chooses to discard all tickets that correspond to other chains, such that if $n$ is large the attacker effectively spends work almost exclusively on the ticket creation. This means that the balance of the two difficulties (ticket creation and block creation) determines how large part of the network CPU power such an attacker would need to control. If ticket creation and block creation require equal work, then the attacker would need to control more than 25% of the whole network. If ticket creation dominates strongly then the fraction needed would approach 50%. Another aspect worth considering is that the multi-strand approach may be better suited for small miners which could be good for decentralization and security.
Implementation of the ticket-based method for a given single-strand blockchain (e.g. Bitcoin or Monero) should be relatively straightforward given that each of the $n$ chains can use the existing software with only slight modifications to include a ticket inside each block, and to add validation of the ticket inside a block when checking if a block is valid.
Depending on the type of information stored, having separate chains may have drawbacks. In case of cryptocurrencies, having multiple chains means that each transaction can only involve outputs that belong to the same chain. So, it would be necessary for users to have a separate “wallet” for each chain. While this might seem awkward, it could potentially be handled by wallet software and the advantage of increased efficiency might be worth that extra complexity, especially if the goal is to handle the transaction rates needed to accommodate daily transactions of billions of users.

### v3

@@
 To avoid the limitation of serially adding one block at a time to a blockchain, multiple semi-independent chains are used such that several valid blocks can be added in parallel, when they are added to separate chains.
-Blocks are added to different chains, the chain index being determined by a “ticket” that the miner must produce before mining a new block.
+Blocks are added to different chains, the chain index being determined by a “ticket” that the miner must produce before creating a new block.
 This allows increasing the transaction rate by several orders of magnitude while the system is still fully decentralized and permissionless, and maintaining security in the sense that a successful attack would require the attacker to control a significant portion of the whole network.
@@
 Information is stored in a distributed ledger, copies of which can exist in many places but no instance has authority over the others.
-The proof-of-work mechanism provides a form of collective decision making regarding how information can be added to the distributed ledger.
+The proof-of-work mechanism provides a form of collective decision-making regarding how information can be added to the distributed ledger.
 One drawback of such blockchain systems is that the rate at which information can be added to the ledger is rather limited.
@@
 For cryptocurrency applications, each block contains one or more transactions, and one way to increase the rate of transactions is to increase the maximum allowed block size.
-However, the block size cannot be increased infinitely; the work involved in added a block needs to be feasible.
+However, the block size cannot be increased infinitely; the work involved in adding a block needs to be feasible.
 In practice, even if the possibilities of increasing block size and reduced block time are taken into account, it appears unlikely that blockchains such as Bitcoin and Monero as currently designed could handle the transaction rates that would be necessary if such cryptocurrencies were to be used by a significant portion of the world’s population.
@@
 The improvement in throughput compared to a single-strand blockchain depends on the chosen amount of parallelism.
-As an example, a multi-stranded chain of 1024 strands would allow 1024 times as many transactions per day.
+As an example, a multi-stranded chain of 1024 strands would in principle allow 1024 times as many transactions per day.
 2 Method
@@
 Instead of a single chain, we consider a different form of distributed ledger that consists of $n=2^{p}$ chains which we label “Chain 0” to “Chain $n-1$”.
-Each chain works similarly to the single-strand case, with the difference that each block is required to include a valid ticket:
-The ticket is a data structure that contains the hash of the last known block from each of the $n$ chains, together with a nonce chosen such that the hash of the ticket satisfies the following: (1) the hash starts with a certain number of zero bits and (2) the last $p$ bits of the hash correspond to the chain index of the chain for which the ticket is valid.
-The procedure to follow for a miner who wants to add a new block now consists of two proof-of-work phases, scanning for first a ticket nonce and then a block nonce, as follows:
-  • First create a ticket by scanning for a ticket nonce that makes the ticket satisfy the criteria (1) and (2) above. Now the miner knows which chain it is possible to add a block to.
-  • Then include that ticket in a new candidate block for the chain that the ticket is valid for, and scan for a block nonce that makes the block valid.
-Note that the procedure above involves two different nonce values, a ticket nonce and a block nonce.
-In both cases the proof-of-work difficulty is determined by the number of zero bits required.
-The difficulties should be selected such that both parts require significant work.
-Note also that the miner cannot choose which chain to add a block to, since the chain index is determined by the ticket.
+Each chain works similarly to the single-strand case, with the difference that each block is required to include a valid ticket and a corresponding signature:
+The ticket is a data structure that contains the hash of the last known block from each of the $n$ chains, together with a pubkey and a nonce chosen such that the hash of the ticket satisfies the following: (1) the hash starts with a certain number of zero bits and (2) the last $p$ bits of the hash correspond to the chain index of the chain for which the ticket is valid.
+The signature in each block is used to provide proof that the block was indeed created by the ticket-holder.
+Before creating a ticket, the miner will first generate a keypair (private key and corresponding pubkey) for the ticket and place the pubkey inside the ticket.
+Later, when a nonce has been found making the ticket valid, the miner creates the block and signs it using the private key before sending the new block to other nodes.
+This ensures that another node cannot “hijack” the ticket and use it to create a different block; only the node that created the ticket has the ability to create the corresponding block.
+Note that there should be a new keypair for each new ticket and the purpose is to allow others to verify the connection between the ticket and the block – the signature shows that whoever created the ticket has also created the corresponding block.
+The procedure to follow for a miner who wants to add a new block now becomes as follows:
+  • Create a keypair to use for the new ticket
+  • Create a ticket by scanning for a ticket nonce that makes the ticket satisfy the criteria (1) and (2) above. Now the miner knows which chain it is possible to add a block to.
+  • Then include that ticket in a new block for the chain that the ticket is valid for, and add a signature proving that the ticket creator has created the block.
+The proof-of-work difficulty is determined by the number of zero bits required. The choice of difficulty can be made in the same way as for the single-strand case.
+Note that the miner cannot choose which chain to add a block to, since the chain index is determined by the ticket.
 The miner tries to find a nonce leading to a hash starting with enough zero bits, while some other bits of the same hash determine the chain index.
@@
   • Check that the required number of zero bits are found in the beginning of the hash of the ticket.
+  • Check that the signature for the block is valid, using the pubkey given inside the ticket.
 The purpose of requiring the ticket to contain the hash of the last block of each chain is to ensure that the ticket is fresh; without that requirement it would be possible for an attacker to spend a long time hoarding tickets to carry out an attack against a specific chain later.
 The requirement to include the hash of the last block of each chain in the ticket makes such hoarding impossible since old tickets will no longer be valid.
+Note that when checking if a ticket is valid, only one of the block hashes inside the ticket is actually checked, namely the one corresponding to the chain index the ticket is for.
+All the other hashes inside the ticket are ignored.
+The reason we need to include all $n$ hashes in the ticket is that when the ticket is created it is not yet known which chain index the ticket will be valid for.
+So we include all the $n$ hashes in the ticket, only one of them will be the relevant one in the end but it is not known which one until later.
+Each of the $n$ chains grows independently of the other chains; while the average time to add a block should be the same, there is nothing restricting the chains to stay at precisely the same length.
+For example, it may happen that Chain 0 is several blocks longer than Chain 1, or the other way around.
 To illustrate how the multi-strand method gives increased efficiency, consider the following example: we take the simplest case $n=2$ so there are two chains, Chain 0 and Chain 1.
@@
 The multi-strand approach discussed in this paper allows increased efficiency because new valid blocks can be added in parallel, which becomes possible when using a set of semi-independent chains instead of a single chain.
+3.1 Sealing blocks using signatures versus using a second proof-of-work phase
+A previous version of this paper suggested to use a second second proof-of-work phase after the ticket creation, as a way to seal the block and protect from the risk of another node hijacking the ticket and using it to create a different block.
+A significant problem with that suggestion is that it is unclear how the difficulty for the second proof-of-work phase would be chosen.
+Miners can have widely varying CPU power, and a choice of difficulty that might be reasonable for one miner would be inadequate for others.
+Therefore, the method presented in Section 2.2 uses only one proof-of-work (for ticket creation) and employs a signature to allow other nodes to verify that the block was created by the ticket creator, thereby establishing the necessary connection between the proof-of-work and the corresponding block.
+3.2 Related work
 The idea of using several chains to increase efficiency is not new; one example of previous work employing this idea is the “Parallel Chains” work of Fitzi et al [3].
@@
 In [3] the miner needs to prepare a metablock effectively containing candidate blocks for each of the $n$ chains, something that is not needed in our ticket-based approach which should therefore have better scalability for large values of $n$.
-The ticket-based method is also simpler in the sense that no special “synchronization chain” is involved, all $n$ chains are treated in the same way, none of them is special and no synchronization between chains is needed.
-Regarding security, the worst case would be an attacker who wants to target one specific chain and chooses to discard all tickets that correspond to other chains, such that if $n$ is large the attacker effectively spends work almost exclusively on the ticket creation.
-This means that the balance of the two difficulties (ticket creation and block creation) determines how large part of the network CPU power such an attacker would need to control.
-If ticket creation and block creation require equal work, then the attacker would need to control more than 25% of the whole network.
-If ticket creation dominates strongly then the fraction needed would approach 50%.
-Another aspect worth considering is that the multi-strand approach may be better suited for small miners which could be good for decentralization and security.
+The ticket-based method is also simpler in the sense that no special “synchronization chain” is involved, all $n$ chains are treated in the same way, none of them is special and no other synchronization between chains is needed beyond the inclusion of last block hashes in the tickets.
+Another related work employing parallel chains is [4] which however differs from our ticket-based method in that [4] requires a miner to create the complete block without knowing which chain the new block will belong to.
+Our ticket-based method employs a two-phase approach where the ticket is created without knowing the chain index, but when the actual block is created the chain index is known, which is important as it allows the miner to select which transactions to include in the block knowing which chain the block will belong to.
+DAG-based protocols such as [5] have some similarity with the present work in the sense that they allow multiple blocks to be generated in parallel.
+However, such methods are much more complicated than the simple scheme we outlined here because they involve complicated relationships between blocks where a block can have multiple predecessors leading to issues such as the “late-predecessor” phenomenon that can affect security and performance [6].
+In our ticket-based method, determining if a new block is valid is much simpler as it only requires access to the chain that block belongs to, nothing more.
+Note that although ticket creation involves hashes of blocks from all chains, validation of a ticket only requires checking one of those hashes, so there is in our case only one predecessor block.
+3.3 Security
+An attacker who wants to target one specific chain would want to create tickets specifically for the targeted chain, but that is not possible since the chain index only becomes known after a ticket is complete.
+So, the attacker is prevented from targeting a specific chain.
+A successful attack becomes possible only if the attacker controls more than half of the whole network, just as in the single-strand case.
+One drawback of the ticket-based multi-strand method as described in Section 2.2 is that a malicious miner who created a valid ticket can then attempt to cause confusion by creating multiple variants of the corresponding block, without redoing the proof-of-work.
+Since the miner has the private key for the ticket it is possible for the miner to sign several different blocks using the same ticket.
+Even though there is nothing to gain in the long run as only one of those blocks can survive, such behavior of malicious miners can increase the likelihood of conflicting versions of the most recently added block for a given chain.
+However, as soon as other blocks are added to the chain the block will be “locked” and protected by the proof-of-work in the added blocks.
+Just as in the single-strand case, for an attacker who controls less than half of the network, the probability of the attacker to catch up diminishes exponentially as more blocks are added.
+Another aspect worth considering is that since the multi-strand approach allows smaller block sizes to be used, it may be better suited for small miners which could be good for decentralization and security.
+3.4 Implementation and usage
 Implementation of the ticket-based method for a given single-strand blockchain (e.g.